# All-in-One Slider for Attribute Manipulation in Diffusion Models

### v1

@@ Introduction
We employ SDXL (Podell et al. 2023) to generate images using the DPM-solver sampler over 50 sampling steps with classifier-free guidance (Ho and Salimans 2022) of 7.5. All the compared methods are implemented following their default configurations available from their official repository. Besides, we edit real images using SDXL-Turbo (Sauer et al. 2024) with the Renoise inversion method (Garibi et al. 2024). We extract features from the 10th and 29th layers of the two text encoders in SDXL, which contain 12 and 32 transformer layers, respectively. We train our Attribute Sparse Autoencoder with 52,000 text samples (52 controllable facial attributes, and 1000 samples per attribute). The details of construction and Attribute Sparse Autoencoder parameters used for the experiments are included in the supplementary material.
Quantitative Metrics
@@
Traditional metrics, such as the CLIP score (Hessel et al. 2021), are limited in their ability to evaluate fine-grained or localized attribute manipulations. These methods typically measure overall semantic similarity, which may obscure whether specific target attributes were correctly edited. To more comprehensively assess image editing quality and manipulation capabilities, we adopt two complementary metrics specifically designed to capture the nuanced effects of facial attribute manipulation. We use the multimodal large language model, Qwen2.5-VL-Instruct (Wang et al. 2024), to evaluate the semantic alignment between the intended textual prompt (e.g., “smile”, “old”, and “makeup”) and the generated image, termed as Qwen Score (QS) (Han et al. 2025). In contrast to traditional metrics, Qwen demonstrates superior reasoning ability and aligns more closely with human judgment. The QS evaluates the edited image relative to the original image based on three criteria: Action Fidelity (the alignment with the edit prompt), Identity Preservation (the consistency of the subject’s identity), and Visual & Anatomical Coherence (the naturalness of the modifications). The average of these three scores is used as the final metric. ID Consistency Score (IS): We further measure the identity preservation between the source image and the edited result using an ID consistency score based on a face recognition model (i.e., Arcface (Deng et al. 2019)). This quantifies how well our method preserves the core identity features despite attribute manipulations. Identity preservation is especially important in real-image editing scenarios, where users expect natural-looking changes without sacrificing personal identity. Higher QS and IS both indicate better editing performance, reflecting more accurate attribute manipulation and stronger identity preservation.
Text-to-Image Attribute Manipulation
To evaluate the ability of All-in-One Slider to manipulate attributes, we present qualitative results in Figure 4. The wide range of visual controls, which covers both semantic attributes (e.g., smile, age, makeup) and physical accessories (e.g., eyeglasses, hat), demonstrates the expressive power and controllability of our method. Each manipulated image preserves the identity and background of the original input, while exhibiting precise and consistent attribute changes. This controllability is enabled by the sparsity and disentanglement of directions in the latent space of Attribute Sparse Autoencoder, where each attribute corresponds to an independent and well-isolated latent direction. This allows precise manipulation without affecting unrelated visual features. In contrast, traditional methods (Baumann et al. 2025; Gandikota et al. 2024) rely on individual sliders per attribute, which inherently restricts the controllable attribute space as each new attribute necessitates its own slider. Our unified framework overcomes this constraint, enabling efficient manipulation across a broader range of attributes.
@@
Multi-Attribute Manipulation
To evaluate the compositional controllability of our method, we perform multi-attribute manipulation experiments as shown in Figure 5. Starting from a base image, we simultaneously apply multiple attributes such as Smile, Makeup, Hat, and Old. The edited images show a visually coherent and semantically consistent blend of attributes without artifacts or conflicts, demonstrating the model’s ability to naturally combine multiple attributes while still preserving the original identity. Similar compositional behavior is also observed in Figure 1 (2), where various attribute combinations (e.g., Smile, Old, Eyeglasses) are applied, yielding realistic and conflict-free results across identities. This compositional behavior stems from our unified latent space of Attribute Sparse Autoencoder, where different attribute directions are effectively disentangled yet compatible. Since all manipulations are represented as sparse vectors within a shared embedding space, their combinations remain semantically coherent and do not interfere with each other.
Quantitative Analysis
We quantitatively evaluate the effectiveness of our method in both single-attribute (Sng) and multi-attribute (Mul) manipulating tasks, using QS and IS as evaluation metrics. We compare against two One-for-One baselines: ConceptSlider (Gandikota et al. 2024) and AttributeControl (Baumann et al. 2025). Results are summarized in Table 1. In the single-attribute setting, our method achieves competitive performance, obtaining the highest IS for the Old attribute and second-best scores for Smile and Makeup in both QS and IS. In contrast, in the multi-attribute setting—where compositional complexity and entanglement pose greater challenges—our method outperforms the baselines. It achieves the highest QS and IS scores across all combinations (Old+Smile, Old+Makeup, and Smile+Makeup). Notably, for the challenging Old+Makeup case, our approach achieves QS of 4.4281 and IS of 0.6277, surpassing both baselines by a clear margin.
@@
Table 1: Quantitative comparisons across single and multiple attributes editing. “QS” denotes the semantic alignment scored by Qwen, and “IS” denotes the identity consistency score. The best and second-best results are highlighted with bold and underline, respectively.
Model Generalization
We further investigate the architectural generalization of our method by applying it to two different diffusion models: Stable Diffusion v1.4 (SD v1.4) and SDXL-Turbo. In Figure 6 (1), progressively increasing the control strength $\lambda$ yields fine-grained variations in makeup intensity, demonstrating effective attribute modulation within the SD v1.4. Figure 6 (2) further validates our method on SDXL-Turbo, where semantic manipulation of Smile attribute is applied with high fidelity and strong identity preservation. These results highlight the robustness and transferability of our approach across distinct models.
Reference-Based Image Editing
To enable high-fidelity editing on real images, we integrate our method with the ReNoise inversion framework (Garibi et al. 2024), which reconstructs latent representations from real inputs. As illustrated in Figure 7, we conduct editing on two representative attributes: Smile (top) and Old (bottom). In the first row, when adding the Smile, our method clearly avoids unnatural facial deformation (e.g., overly rounded cheeks for AttControl). In the second case, it successfully preserves accessories like eyeglasses during the addition of complex attributes such as Christmas elements with the addition of old. Compared to AttributeControl, our approach better maintains critical identity features, which is primarily attributed to our Attribute Sparse Autoencoder-based decoupling strategy, which minimizes unintended attribute entanglement and improves the overall stability and locality of edits via sparse semantic modulation.
Figure 6: Qualitative results on SD v1.4 (1) and SDXL-Turbo (2) showing the effectiveness of our method in attribute modulation (e.g., makeup intensity and smile generation) across U-Net-based architectures.
@@
Figure 7: Real image editing. We apply our method to edit Smile (top) and Old (bottom) attributes using the ReNoise inversion framework (Garibi et al. 2024). Compared to AttributeControl, our method better preserves identity details like eyeglasses and facial structure.
@@
Figure 8: Continuous zero-shot generalization to different racial attributes. The model progressively manipulates facial features across varying strengths of the racial attribute ($\lambda=0.3,0.4,0.5$).
Zero-shot Generalization Ability
@@
Although our All-in-One Slider is trained on a fixed set of facial attributes, it demonstrates strong zero-shot generalization to new racial and identity attributes not encountered during training. As shown in Figure 8, the model can continuously modify facial attributes, including races (e.g., African, Korean, Chinese, Hispanic, and Indian) and celebrity-like identities (e.g., Barack Obama and Einstein in Figure 1 (3)), without the need for additional fine-tuning or supervision.
@@
Ablation Study
@@
Layer Selection
We evaluate manipulating performance across different combinations of transformer layers, spanning from intermediate (8/28) to deep (10/28) configurations in SDXL’s dual text encoders. These layers have the most significant impact on generated images, covering both shallow and deep semantic regions. As shown in Table 2, deeper layers generally yield higher Qwen scores (semantic score), while identity scores tend to decline. The 10/28 combination achieves the best overall balance, indicating that mid-level layers are more suitable for simultaneously preserving semantic intent and identity. This clearly reflects the practical importance of choosing appropriate feature granularity for localized and identity-sensitive editing in practice.
Effect of Manipulating Strength $\lambda$
Figure 9 shows how varying $\lambda$ affects the trade-off between manipulation strength and identity retention. Lower values (e.g., $\lambda=0.15$) lead to under-editing, whereas higher values (e.g., $\lambda=0.30$) result in stronger attribute expression. We observe that $\lambda=0.20$ achieves a favorable balance, yielding strong attribute alignment (QS) while maintaining high IS across all attributes. In particular, degradation is most severe on “old”, suggesting that aging transformations are more entangled with identity features. These results demonstrate the inherent antagonistic effect between attribute change and identity preservation in the attribute manipulation.
@@
Table 2: Evaluation for attributes old, smile, and makeup across different Manipulation Layer (X/Y) of SDXL’s dual text encoders, where X is the selected layer index in the first encoder and Y in the second.
Figure 9: Effect of manipulating strength ($\lambda$) on attribute manipulation (Qwen score) and identity preservation (ID consistency) for different attributes (old, smile, makeup).
Conclusion
We propose the All-in-One Slider, a sparse and lightweight framework for controllable face image generation using the Attribute Sparse Autoencoder. Unlike previous approaches that rely on attribute-specific training, our approach introduces a unified and disentangled latent space, enabling fine-grained, continuous manipulation of multiple facial attributes while preserving identity and visual fidelity. Extensive experiments demonstrate strong performance of our method in both single- and compositional-attribute manipulation, as well as zero-shot generalization to unseen attributes. All-in-One Slider presents a new and unified paradigm for continuous attribute control of image generation, offering scalable, interpretable, and flexible advantages. It is promising to inspire more work to achieve continuous attribute control with low training and parameter cost based on the unified space.
Supplementary Material
1    Implementation Details
Training Details
We use SDXL as the base generative model and extract intermediate text embeddings from its 11th (of 12) and 29th (of 32) transformer layers blocks (layer.10 and layer.28) across the dual encoders to train the Attribute Sparse Autoencoder. The activations are computed over prompts covering 52 attributes, each augmented with diverse identity-aware prefixes to encourage broad semantic coverage. The dimension of text embedding is 2048 (concatenation of 768 and 1280). The autoencoder has a latent dimension of 32,768 with an expansion factor of 16. It is trained using the Adam optimizer with a learning rate of $4\times 10^{-4}$ and a batch size of 4096 latent tokens per step. The total training budget is 400 million tokens, corresponding to approximately 97,656 training steps. During attribute selection, we retain the top-$k=128$ most activated dimensions. To mitigate overfitting and encourage broader attribute coverage, we introduce an auxiliary selection mechanism with $k_{\text{aux}}=256$ and apply a regularization coefficient of $\alpha=0.1$ to increase its influence during training. All experiments were performed using a single NVIDIA GeForce RTX 4090 GPU with 24 GB of VRAM.
Training Prompt Construction
@@ 1    Implementation Details
To enrich visual diversity and reduce identity bias, we prepend each prompt with descriptive prefixes drawn from a fixed pool. These include phrases like “a photo of a”, “a photo of a beautiful”, as well as identity-related descriptors such as “a photo of a black/white/asian/arab/caucasian/african-american”, and their compositional forms (e.g., “a photo of a beautiful asian”).
@@
Image Editing Evaluation Criteria
In the evaluation of image edits, we introduce Qwen Score (QS) to assess changes in motion or facial expression. This is done using a 5-point scale across three dimensions: Action / Expression Fidelity, Identity Preservation, and Visual & Anatomical Coherence. This system is based on the scoring framework provided by the Qwen2.5-VL-Instruct model. The criteria are as follows:
Expression Fidelity captures how faithfully the edited image reflects the target expression described in the instruction. A high score reflects a precise match in terms of gesture, intensity, and orientation. In contrast, lower scores are triggered by absent or incorrectly rendered expressions, partial changes (e.g., only some facial muscles involved), or misaligned pose dynamics.
@@
Each image pair (before and after editing) is evaluated in terms of three dimensions: Action Fidelity, Identity Preservation, and Visual & Anatomical Coherence, each scored on a 5-point scale. In addition to the numerical scores, a brief reasoning (within 20 words) is provided to justify the evaluation results. As an example, we present the evaluation of an edited image, for which the Qwen model assigned scores of 4 for Action Fidelity, 5 for Identity Preservation, and 4 for Visual & Anatomical Coherence (see Figure 10). The edited image shows a successful transition from a neutral expression to a clear, natural-looking smile, while preserving the head pose and facial features. Minor inaccuracies in the smile’s angle or intensity lead to an Action Fidelity score of 4. The subject’s identity is perfectly preserved, justifying a score of 5 for Identity Preservation. The image also maintains visual and anatomical coherence, with only slight imperfections, resulting in a score of 4 for Visual & Anatomical Coherence. According to Qwen’s analysis, the brief explanation for this result is: “The subject’s expression has changed to a smile, but the overall pose remains the same”.
@@
2    Additional Experimental Results
Attribute Manipulation

### v2

@@ 4.1 Experimental Setup
Implementation Details. We employ SDXL [34] to generate images using the DPM-solver sampler over 50 sampling steps with classifier-free guidance [19] of 7.5. All the compared methods are implemented following their default configurations available from their official repository. Besides, we edit real images using SDXL-Turbo [39] with the Renoise inversion method [15]. We extract features from the 11th and 29th layers (layer.10 and layer.28) of the two text encoders in SDXL, which contain 12 and 32 transformer layers, respectively. We train our Attribute Sparse Autoencoder with 52,000 text samples (52 controllable facial attributes, and 1000 samples per attribute). The details of construction and Attribute Sparse Autoencoder parameters used for the experiments are included in the supplementary material.
@@
Quantitative Metrics. Traditional metrics, such as the CLIP score [18], are limited in their ability to evaluate fine-grained or localized attribute manipulations. These methods typically measure overall semantic similarity, which may obscure whether specific target attributes were correctly edited. To more comprehensively assess image editing quality and manipulation capabilities, we adopt two complementary metrics specifically designed to capture the nuanced effects of facial attribute manipulation. We use the multimodal large language model, Qwen2.5-VL-Instruct [43], to evaluate the semantic alignment between the intended textual prompt (e.g., “smile”, “old”, and “makeup”) and the generated image, termed as Qwen Score (QS) [16]. In contrast to traditional metrics, Qwen demonstrates superior reasoning ability and aligns more closely with human judgment. The QS evaluates the edited image relative to the original image based on three criteria: Action Fidelity (the alignment with the edit prompt), Identity Preservation (the consistency of the subject’s identity), and Visual & Anatomical Coherence (the naturalness of the modifications). The average of these three scores is used as the final metric. Moreover, we further measure the identity preservation between the source image and the steered result using an ID Consistency Score (IS) based on a face recognition model (i.e., Arcface [11]). This quantifies how well our method preserves the core identity features despite attribute manipulations. Identity preservation is especially important in real-image editing scenarios, where users expect natural-looking changes without sacrificing personal identity. Higher QS and IS both indicate better editing performance, reflecting more accurate attribute manipulation and stronger identity preservation.
@@
4.2 Text-to-Image Attribute Manipulation
To evaluate the ability of All-in-One Slider to manipulate attributes, we present qualitative results in Figure 4. The wide range of visual controls, which covers both semantic attributes (e.g., smile, age, makeup) and physical accessories (e.g., eyeglasses, hat), demonstrates the expressive power and controllability of our method. Each manipulated image preserves the identity and background of the original input, while exhibiting precise and consistent attribute changes. This controllability is enabled by the sparsity and disentanglement of directions in the latent space of Attribute Sparse Autoencoder, where each attribute corresponds to an independent and well-isolated latent direction. This allows precise manipulation without affecting unrelated visual features. In contrast, existing methods [1, 13] rely on individual sliders per attribute, which inherently restricts the controllable space as each new attribute requires its own slider. Our unified framework overcomes this constraint, enabling efficient manipulation across broader attributes.
@@ 4.2 Text-to-Image Attribute Manipulation
Multi-Attribute Manipulation. To evaluate the compositional controllability of our method, we perform multi-attribute manipulation experiments as shown in Figure 5. Starting from a base image, we simultaneously apply multiple attributes such as Smile, Makeup, Hat, and Old. The edited images show a visually coherent and semantically consistent blend of attributes without artifacts or conflicts, demonstrating the model’s ability to naturally combine multiple attributes while still preserving the original identity. Similar compositional behavior is also observed in Figure 1 (2), where various attribute combinations (e.g., Smile, Old, Eyeglasses, WavyHair) are applied, yielding realistic and conflict-free results across identities. This compositional behavior stems from our unified latent space of Attribute Sparse Autoencoder, where different attribute directions are effectively disentangled yet compatible. Since all manipulations are represented as sparse vectors within a shared embedding space, their combinations remain semantically coherent and do not interfere with each other.
Table 1: Quantitative comparisons of single and multiple attributes editing. “QS” denotes the semantic alignment score by Qwen, and “IS” denotes the identity consistency score. The best and second-best results are highlighted with bold and underline, respectively.
@@
Quantitative Analysis. We quantitatively evaluate the effectiveness of our method in both single-attribute (Sng) and multi-attribute (Mul) manipulating tasks, using QS and IS as evaluation metrics. We compare against two One-for-One baselines: ConceptSlider [13] and AttributeControl [1]. Results are summarized in Table 1. In the single-attribute setting, our method achieves competitive performance, obtaining the highest IS for both Old and Makeup attributes, and ranking first in QS for the Old attribute and second for Smile and Makeup. In contrast, in the multi-attribute setting—where compositional complexity and entanglement pose greater challenges—our method outperforms the baselines. It achieves the highest QS across all combinations (Old+Smile, Old+Makeup, and Smile+Makeup) and the highest IS scores in two out of three cases. Notably, for the challenging Old+Makeup case, our approach achieves a QS of 4.4281, surpassing the second-best baseline (4.0555) by a clear margin while still achieving a highly competitive IS of 0.6277.
@@
Figure 6: Real image editing. We apply our method to edit Smile (top) and Old (bottom) attributes using the ReNoise inversion framework [15]. Compared to AttControl, our method better preserves identity details like eyeglasses and facial structure.
4.3 Reference-Based Image Editing
To enable high-fidelity editing on real images, we integrate our method with the ReNoise inversion framework [15], which reconstructs latent representations from real inputs. As illustrated in Figure 6, we conduct editing on two representative attributes: Smile (top) and Old (bottom). In the first row, when adding the Smile, our method clearly avoids unnatural facial deformation (e.g., overly rounded cheeks observed in AttControl). In the second case, it successfully preserves accessories like eyeglasses during the addition of complex attributes, such as combining Christmas elements with the old attribute. Compared to AttributeControl, our approach better maintains critical identity features. This is primarily attributed to our Attribute Sparse Autoencoder-based decoupling strategy, which minimizes unintended attribute entanglement and improves the overall stability and locality of edits via sparse semantic modulation.
@@ 4.3 Reference-Based Image Editing
Figure 7: Continuous zero-shot generalization to different racial attributes. The model progressively manipulates facial features across varying strengths of the racial attribute.
4.4 Zero-shot Generalization Ability
@@
Although our All-in-One Slider is trained on a fixed set of facial attributes, it demonstrates strong zero-shot generalization to new racial and identity attributes not encountered during training. As shown in Figure 7, the model can continuously modify facial attributes, including races (e.g., African, Chinese and Indian) and celebrity-like identities (e.g., Barack Obama and Einstein in Figure 1 (3)), without the need for additional fine-tuning or supervision.
@@
4.5 Applications
Model Generalization. We further investigate the generalization of our method by applying it to various diffusion models, including Stable Diffusion v1.4 (SD v1.4), SDXL-Turbo and the Diffusion Transformer (DiT) (see Appendix). In Figure 8 (1), progressively increasing the control strength $\lambda$ yields fine-grained variations in makeup intensity, demonstrating effective attribute modulation within the SD v1.4. Figure 8 (2) further validates our method on SDXL-Turbo, where semantic manipulation of the Smile attribute is achieved with high fidelity and robust identity preservation. These results highlight the robustness and transferability of our approach across diverse backbones.
Figure 8: Qualitative results on SD v1.4 (1) and SDXL-Turbo (2) showing the effectiveness of our method in attribute modulation (e.g., makeup and smile) across U-Net-based architectures.
Style Manipulation. To validate our method’s versatility beyond facial attributes manipulation, we apply it to photography style steering. We train SAE on 40 different photography styles following the procedure in Sec. 3.2, including Black-and-White, Golden-hour, and Fisheye, etc. As shown in Figure 18, our method effectively transforms the original image into various photography styles while preserving the core content and composition.
Figure 9: Style manipulation on SDXL: the original image is progressively transformed into Black-and-White and Golden-hour styles while preserving main content and composition.
Multi-Subject Attribute Manipulation. To extend the universal Attribute Sparse Autoencoder (detailed in Sec. 3.2) to multi-subject scenarios, we introduce an Attention Pooling (AttPooling) Aggregator (AAg) module. It aggregates the latent features of attribute semantics ($z^{\pm}$) onto a token, enabling precise localization of the target subject (e.g., “woman” or “man”). To extract pure attribute directions in this setting, we create 4,000 sentence pairs, with each pair consisting of one sentence describing a specific attribute (e.g., smile, old, makeup) and another without it. Building upon the pretrained universal SAE (i.e., trained in stage 1 in Sec. 3.2), we jointly fine-tune the SAE and the AAg module with SDXL using the following objective: $\mathcal{L}_{\text{multi}}=\mathcal{L}_{\text{sae}}+\eta\mathcal{L}_{\text{cons}}$. Here, $\mathcal{L}_{\text{sae}}$ denotes the original sparse and reconstruction loss (Eq. 5), while $\mathcal{L}_{\text{cons}}$ is a consistency loss that encourages the model to preserve non-target regions. This fine-tuning process enhances the model’s ability to apply attributes selectively to specified subjects. The overall fine-tuning framework for multi-subject manipulation is illustrated in Figure 10(a) (detailed in Appendix). As shown in Figure 10(b), our method can effectively manipulate attributes for specific subjects (e.g., adding old to the man, and smile to the woman) while maintaining the identity and context of other elements in the scene.
Figure 10: (a) Framework for fine-tuning multi-subject manipulation. The Attribute Sparse Autoencoder is fine-tuned using paired sentences combined the diffusion model, and we introduce an Attpooling Aggregator (AAg) module to locate the target subject for manipulation. (b) Qualitative results of applying attributes to the targeted subject (e.g., old to the man, smile to the woman).
4.6 Ablation Study
@@
Layer Selection. We evaluate performance across various layer configurations of SDXL’s dual text encoders, specifically testing different combinations of layer indices (X/Y) as shown in Table 2. These layers exert a pronounced impact on the generated images, capturing different levels of semantic representation. The results indicate that deeper layers generally yield higher QS (semantic score), while IS tend to decline. The 10/28 combination achieves the best overall balance, indicating that mid-level layers are more effective at simultaneously preserving semantic intent and identity. This clearly reflects the practical importance of choosing appropriate feature granularity for localized and identity-sensitive editing in practice.
Effect of Manipulating Strength $\lambda$. Figure 11 illustrates how varying $\lambda$ affects the trade-off between attribute control and identity preservation. Lower values (e.g., $\lambda=0.15$) lead to under-editing, whereas higher values (e.g., $\lambda=0.30$) result in stronger attribute expression. In particular, semantic degradation is most pronounced for the old, suggesting that aging transformations are more deeply entangled with identity features. These results demonstrate the inherent trade-off between attribute modification and identity preservation during attribute manipulation.
Table 2: Evaluation for attributes across different Manipulation Layer (X/Y) of SDXL’s dual text encoders, where X is the selected layer index in the first encoder and Y in the second.
@@
Figure 11: Effect of manipulating strength $\lambda$ on attribute expression (QS) and identity preservation (IS) for various attributes (old, smile, makeup).
5 Conclusion
We propose the All-in-One Slider, a sparse and lightweight framework for controllable attribute manipulation using the Attribute Sparse Autoencoder. Unlike prior approaches that rely on attribute-specific training, our method introduces a unified and disentangled latent space, enabling fine-grained, continuous manipulation of multiple facial attributes while preserving identity and visual fidelity. Extensive experiments demonstrate performance of our method in both single- and compositional-attribute manipulation, as well as zero-shot generalization to unseen attributes. All-in-One Slider presents a new and unified paradigm for continuous attribute control of image generation, offering scalable, interpretable, and flexible advantages. It is promising to inspire more work to achieve continuous attribute control with low training and parameter costs based on the unified space.
Acknowledgments. This work has been supported by the Fundamental Research Funds for the Central Universities (No. 2022XKRC015), Natural Science Foundation of Beijing Municipality (No. L252025), National Natural Science Foundation of China (No. 62372033), and the Science and Technology Development Fund of Macao SAR under the International Collaborative Research (No. 0001/2025/AIJ).
Appendix A Implementation Details
Training Details.
We use SDXL as the base generative model and extract intermediate text embeddings from its 11th (of 12) and 29th (of 32) transformer layers blocks (layer.10 and layer.28) across the dual encoders to train the Attribute Sparse Autoencoder. The activations are computed over prompts covering 52 attributes, each augmented with diverse identity-aware prefixes to encourage broad semantic coverage. The dimension of text embedding is 2048 (concatenation of 768 and 1280). The autoencoder has a latent dimension of 32,768 with an expansion factor of 16. It is trained using the Adam optimizer with a learning rate of $4\times 10^{-4}$ and a batch size of 4096 latent tokens per step. The total training budget is 400 million tokens, corresponding to approximately 97,656 training steps. During attribute selection, we retain the top-$k=128$ most activated dimensions. To mitigate overfitting and encourage broader attribute coverage, we introduce an auxiliary selection mechanism with $k_{\text{aux}}=256$ and apply a regularization coefficient of $\alpha=0.1$ to increase its influence during training. All experiments were performed using a single NVIDIA GeForce RTX 4090 GPU.
Training Prompt Construction.
@@ Appendix A Implementation Details
To enrich visual diversity and reduce identity bias, we prepend each prompt with descriptive prefixes drawn from a fixed pool. These include phrases like “a photo of a”, “a photo of a beautiful”, as well as identity-related descriptors such as “a photo of a black / white / asian / arab / caucasian”, and their compositional forms (e.g., “a photo of a beautiful asian”).
@@
Image Editing Evaluation Criteria.
In the evaluation of image edits, we introduce Qwen Score (QS) to assess semantic attribute changes or facial expression. This is done using a 5-point scale across three dimensions: Action / Expression Fidelity, Identity Preservation, and Visual & Anatomical Coherence. This system is based on the scoring framework provided by the Qwen2.5-VL-Instruct model. The criteria are as follows:
Expression Fidelity captures how faithfully the edited image reflects the target expression described in the instruction. A high score reflects a precise match in terms of intensity and orientation. In contrast, lower scores are triggered by absent or incorrectly rendered expressions, partial changes (e.g., only some facial muscles involved), or misaligned pose dynamics.
@@
Each image pair (before and after steering) is evaluated in terms of three dimensions: Expression Fidelity, Identity Preservation, and Visual & Anatomical Coherence, each scored on a 5-point scale. In addition to the numerical scores, a brief reasoning (within 20 words) is provided to justify the evaluation results. As an example, we present the evaluation of an manipulated image, for which the Qwen model assigned scores of 4 for Action Fidelity, 5 for Identity Preservation, and 4 for Visual & Anatomical Coherence (see Figure 12). The manipulated image shows a successful transition from a neutral expression to a clear, natural-looking smile, while preserving consistency in the head pose and facial features. Minor inaccuracies in the smile’s angle or intensity lead to an Action Fidelity score of 4. The subject’s identity is perfectly preserved, justifying a score of 5 for Identity Preservation. The image also maintains visual and anatomical coherence, with only slight imperfections, resulting in a score of 4 for Visual & Anatomical Coherence. According to Qwen’s analysis, the brief explanation for this result is: “The subject’s expression has changed to a smile, but the overall pose remains the same”.
@@
Multi-Subject Manipulation.
The goal of multi-subject manipulation is to steer the facial attributes of a specific target subject (e.g., a man or a woman) within a multi-subject scene, while preserving the other subjects and the rest of the scene consistency. As shown in Sec.4.5 in the main paper, we employ the fine-tuning objective $\mathcal{L}_{\text{multi}}=\mathcal{L}_{\text{sae}}+\eta\mathcal{L}_{\text{cons}}$, where $\eta$ is set to 0.1. During inference, we first encode the paired sentences hidden states $\text{H}^{+}$ and $\text{H}^{-}$ (with and without the target attribute) into their corresponding sparse latent representations, denoted as $z^{+}$ and $z^{-}$, using the pre-trained Attribute Sparse Autoencoder. These representations are then passed through the Attention Pooling (AttPooling) Aggregator (AAg) module, which aggregates the attribute-related semantics into compact token-level embeddings. The aggregated directional vector, $\Delta z=\text{AAg}(z^{+})-\text{AAg}(z^{-})$, captures the pure attribute semantics. This vector is then decoded via the SAE decoder and added to the target subject’s embedding $e_{\text{target}}$, yielding the updated subject embedding: $e_{\text{target}}^{\prime}=e_{\text{target}}+\text{DEC}(\lambda\cdot\Delta z)$. This process ensures that the attribute manipulation is precisely localized to the target subject, while preserving the identity and context of other elements in the scene.
Photography Style List.
We train our model on a diverse set of photography styles, including: architectural, astrophotography, black-and-white, blue hour, bokeh, cinematic, documentary, dreamy, fashion, film grain, fisheye, food, golden hour, HDR, high contrast, infrared, landscape, long exposure, low light, macro, minimalist composition, monochrome, muted colors, natural lighting, neon lighting, overexposed, pastel tone, portrait, product, sepia tone, soft focus, street photography, studio lighting, surreal, telephoto, underexposed, vintage, wide-angle.
Appendix B Additional Experimental Results
B.1 Additional attribute Manipulation